OpenStudy (shaniehh):

Quick help What is the value of 373 using the identity (x − y)3 = x3 − 3x2y + 3xy2 − y3? Hint: 373 = (40 − 3)3; therefore, x = 40 and y = 3.

OpenStudy (shaniehh):

Please explain

OpenStudy (monkeyagf):

what grade is that?

OpenStudy (shaniehh):

What is the value of 37^3 using the identity (x − y)^3 = x^3 − 3x2y + 3xy^2 − y^3? Hint: 37^3 = (40 − 3)^3; therefore, x = 40 and y = 3.